# Generalized Animal Imitator: Agile Locomotion with Versatile Motion Prior

### v2

@@ Abstract
The agility of animals, particularly in complex activities such as running, turning, jumping, and backflipping, stands as an exemplar for robotic system design. Transferring this suite of behaviors to legged robotic systems introduces essential inquiries: How can a robot be trained to learn multiple locomotion behaviors simultaneously? How can the robot execute these tasks with a smooth transition? How to integrate these skills for wide-range applications? This paper introduces the Versatile Instructable Motion prior (VIM) – a Reinforcement Learning framework designed to incorporate a range of agile locomotion tasks suitable for advanced robotic applications. Our framework enables legged robots to learn diverse agile low-level skills by imitating animal motions and manually designed motions. Our Functionality reward guides the robot’s ability to adopt varied skills, and our Stylization reward ensures that robot motions align with reference motions. Our evaluations of the VIM framework span both simulation environments and real-world deployment. To the best of our knowledge, this is the first work that allows a robot to concurrently learn diverse agile locomotion skills using a single learning-based controller in the real world. Further details and supportive media can be found at our project site: https://rchalyang.github.io/VIM
@@
Figure 1: Real-Robot Trajectory. Our robot demonstrates diverse agile locomotion skills, including running, jumping, and back-flipping in diverse real-world scenarios using a single motion prior and without any fine-tuning.
*Equal contribution
I Introduction
@@
Researchers have been studying for years equipping legged robots with agility comparable to that of natural quadrupeds. Picture a golden retriever gracefully maneuvering in a park: darting, leaping over obstacles, and pursuing a thrown ball. These tasks, effortlessly performed by many animals, remain challenging for contemporary legged robots. To accomplish such tasks, robots need not only master individual agile locomotion skills like running and jumping but also the capacity to adaptively select and configure these skills based on sensory inputs. We regard this kind of complicated task requiring highly agile locomotion skills as advanced parkour for legged robots. The inherent ability of quadrupeds to smoothly execute diverse locomotion skills across varied tasks inspires our pursuit of a control system with a general locomotion motion prior that includes these skills. In this direction, we introduce a novel RL framework, Versatile Instructable Motion prior (VIM) aiming to endow legged robots with a spectrum of reusable agile locomotion skills by integrating existing agile locomotion knowledge.
@@
Historically, agile gaits[4, 36, 37] for legged robots have been sculpted using model-based or optimization methods. While promising, these methods demand significant engineering input and precise state estimation. Learning-based controllers enable robots to walk or run while addressing these limitations, although they still fall behind in agility. Imitation-based controllers are also proposed to learn from motion sequences from animals [40] or optimization methods [16]. Research on incorporating visual observations [1, 57, 56, 24, 60, 27, 62] or elevation maps [33, 44] further enables legged robots to traverse complex terrain like stones. Despite encouraging results, most of these works focus on building a single controller from scratch, even though much of the learned locomotion skills could be shared across tasks. Recent works start building a reconfigurable low-level motion prior [46, 23, 5, 42, 26, 22] for downstream applications. However, previous methods failed to make the best use of existing skills to learn diverse locomotion skills with high agility.
@@
Figure 2: Learning Agile Locomotion Skills from Reference Motions: Our system learns a single instructable motion prior, from a diverse reference motion dataset and makes the learned skills available for solving downstream tasks.
@@
In this work, we focus on building low-level motion prior to utilize existing locomotion skills in nature and previous optimization methods, and learn multiple highly agile locomotion skills simultaneously, as in Figure 2. Even though we cannot fully comprehend the agility of animals and lack a unified framework for model-based controls, we recognize that motion sequences offer a consistent representation of diverse agile locomotion skills. Our motion prior extracts and assimilates a range of locomotion skills from reference motions, effectively mirroring their dynamics. These references comprise motion capture (mocap) sequences from quadrupeds, synthetic motion sequences complementing mocap data, and optimized motion trajectories. Throughout the training phase, we translate varied reference motion clips into a unified latent command space, guiding the motion prior to recreate locomotion dynamics based on these latent commands and the robot’s inherent state.
For legged robots, we define a locomotion skill as the ability of the robot to produce a specific trajectory. To break down the intricacies of movement, we classify it into two primary aspects: Functionality and Style. Functionality pertains to the fundamental movement objectives a robot aims to achieve, such as advancing forward at a predefined speed. Style, in contrast, focuses on the specific mechanics of how a robot accomplishes a task. To illustrate, two robots might be programmed to progress at an identical speed, but the intricacies of their movement—like step size or frequency—might differ considerably. Simultaneously instructing a robot in both these domains is nontrivial [55]. Drawing inspiration from how humans learn complicated tasks, especially in fields demanding physical prowess like athletics, we identify three core feedback modalities: objective performance metrics, qualitative assessments, and detailed kinematic guidance. Adopting this structured feedback approach, our robot starts with mastering the basic functional objective and subsequently turns into refining the detailed locomotion gaits.
@@
By incorporating diverse reference motions and our reward design, our Versatile Instructable Motion prior (VIM) learns diverse agile locomotion skills and makes them available for intricate downstream tasks. With our VIM, we enable legged robots to perform advanced robotics parkour in the real world. We also evaluate our method in the simulation and real world, as Figure 1. Our method significantly outperforms baselines in terms of final performance and sample efficiency.
II Related Work
@@
Blind Legged Locomotion: Classical legged locomotion controllers [19, 59, 52, 34, 4, 43] based on model-based methods [8, 17, 10, 12, 3, 20, 49] and trajectory optimization [7, 37] have shown promising results in diverse tasks with high levels of agility. Nonetheless, these methods normally come with considerable engineering work for the specific task, high computation requirements during deployment, or fragility to complex dynamics. Learning-based methods [1, 29, 33, 47, 9, 15, 31] controllers are proposed to offer robust and lightweight controllers for deployment at the cost of offline computation. Peng et al [14] developed a controller producing non-agile life-like gaits by imitating animals. Though previous works offer robust or agile locomotion controllers across complex environments, these works focus on finishing a single task at a time without reusing previous experience. Peng et al [41] leverage reference motions as prior knowledge when directly addressing specific tasks. Li et al [30] obtain an agile locomotion skill from a single partial rough demonstration including only robot root trajectory. Smith et al. [48] utilize existing locomotion skills to solve specific downstream tasks. Vollenweider et al. [54] utilize multiple AMP [41] to develop a controller to solve a fixed task set. In this paper, our motion prior captures diverse agile locomotion skills from reference motions generated by trajectory optimization and provides them for intricate future downstream tasks.
@@
Motion Priors: Due to the low sample efficiency and considerable effort required for reward engineering of RL, low-level skill pre-training has drawn growing attention in recent years. Singh et al [46] utilize a flow-based model to build an actionable motion prior with motion sequences generated by scripts. More recent works [23, 5, 42, 26, 13, 22] focus on building low-level motion prior for downstream tasks but fail to include diverse highly agile locomotion skills. Peng et al [42] develop a simulation-based low-level motion prior entirely through unsupervised methods, yet they do not assure the acquisition of agile skills from the reference motion dataset. In this work, we build motion prior with reference motions consisting of mocap sequences, synthesized motion sequences, and trajectories from optimization methods and learn multiple highly agile locomotion skills with a single controller.
III Learn Versatile Instructable Motion Prior
We present the Versatile Instructable Motion prior (VIM), depicted in Figure 3, designed to acquire a wide range of agile locomotion skills concurrently from multiple reference motions. The development of our motion prior involves three stages: assembling a comprehensive dataset of reference motions sourced from diverse origins, crafting a motion prior that processes varying reference motions and the robot’s proprioceptive feedback to generate motor commands, and finally, utilizing an imitation-based reward mechanism to effectively train this motion prior.
III-A Reference motion dataset
TABLE I: Reference Motion Dataset: Our reference motions (11 reference motions in total) come from motion capture of animal motion, trajectory optimization method, and synthesized data with generative model whose length ranges from 32 to 500. During training, we repeat the reference motions cyclically to fit the length of the episode.
Motion length 500 32 500 38 45 38
Motion length 500 500 129 120 64
Our primary objective is to curate a skill set for the robot that covers diverse functions and agility levels, equipping it to handle complex downstream tasks. Our dataset encompasses reference motions for locomotion skills, including but not limited to canter, pace, walk, trot, turns, backflips, and various jumps. These reference motions are derived from: (a) Mocap data of quadrupeds, specifically a subset from previous work [61], despite its inherent challenges like noise due to the unpredictability of animal behavior; (b) synthesized motions generated using a generative model [61], aimed at enhancing dataset diversity by capturing challenging locomotion actions; (c) motions crafted through trajectory optimization methods.
@@ III-A Reference motion dataset
To address the methodology disparities between the real animals and our robot, we retarget both mocap and synthesized sequences to our robot as Peng et al. [40]. While mocap and synthesized motions offer extensive data, not all sequences may be practically achievable by the robot. Thus, our dataset is supplemented with motion sequences from trajectory optimization, emphasizing intricate moves like jumps and backflips. The comprehensive reference motion list can be found in Table I. Each trajectory in our dataset, represented as $(s^{\text{ref}}_{0},s^{\text{ref}}_{1},\cdots,s^{\text{ref}}_{T})$, where $s^{\text{ref}}_{i}$ is the robot state at $i$th timestep in the reference motion, and $T$ is the length of the reference motion. For every frame of the reference motion, it includes the following information: 1) robot pose in the world frame 2) robot velocity in the robot frame 3) foot position in the robot frame 4) foot height over the ground 5) robot joint angle, and 6) robot joint velocity. Note that reference motions do not include specific motor commands which are absent in the captured and synthesized data. For information that is hard to obtain in the real world, like robot pose and robot velocity, we only use them in simulation for reward calculation during training, our policy does not require that information in the real world. We denote the dataset as $\mathcal{D}=\{(s^{\text{ref}}_{0},s^{\text{ref}}_{1},\cdots,s^{\text{ref}}_{T})_{i}\}_{i=1}^{N}$, where $N$ is the total number of reference motions we have in the dataset.
III-B Motion Prior Structure
@@
Figure 3: Versatile Instructable Motion prior (VIM): Our instructable motion prior consists of a reference motion encoder and a low-level policy. Our reference motion encoder maps reference motions into latent skill space indicating target robot pose and low-level policy output motor command. $V_{low}$ is the low-level critic for RL training. Reward Design: Our reward includes Functionality reward and Style reward. Our Functionality reward encourages the robot to track the root pose trajectory of the reference motion, and Style reward encourages the robot to track the joint motion of the reference motion.
@@
Reference motion encoder: Our reference motion encoder $\mathbb{E}_{\text{ref}}(\cdot)$ maps segments of reference motion to latent commands in a latent skill space that outlines the robot’s prospective movement. These segments are expressed as $\hat{s}^{\text{ref}}_{t}=\{s^{\text{ref}}_{t+1},s^{\text{ref}}_{t+2},s^{\text{ref}}_{t+5},s^{\text{ref}}_{t+10},s^{\text{ref}}_{t+30}\}$. Specifically, we choose $s^{\text{ref}}_{t+1},s^{\text{ref}}_{t+2},s^{\text{ref}}_{t+5}$ to provide immediate desired future motion, $s^{\text{ref}}_{t+10},s^{\text{ref}}_{t+30}$ to provide desired motion over a longer time-span. We model the latent command as a Gaussian distribution $\mathcal{N}(\mathbb{E}_{\text{ref}}^{\mu}(\hat{s}^{\text{ref}}_{t}),\mathbb{E}_{\text{ref}}^{\sigma}(\hat{s}^{\text{ref}}_{t}))$ from which we draw a sample at each interval to guide the low-level policy.
To maintain a temporal-consistent latent skill space, our training integrates an information bottleneck [51, 2] objective $L_{\text{AR}}$, where the prior follows an auto-regressive model [6]. Specifically, given the sampled latent command for the previous time step $z_{t-1}$, we minimize the KL divergence between the current latent Gaussian distribution and a Gaussian prior parameterized by $z_{t-1}$,
$L_{\text{AR}}(\hat{s}^{\text{ref}}_{t},z_{t-1})=\beta\text{KL}\left(\mathcal{N}(\mu_{t},\sigma_{t}^{2})\parallel\mathcal{N}(\alpha z_{t-1},(1-\alpha^{2})I)\right),$
where $\alpha=0.95$ is the scalar controlling the effect of correlation, $\beta$ is the coefficient balancing regularization.
Low-level policy training: Our low-level policy $\pi_{\text{low}}$ takes latent command $z_{t}$ representing the desired robot pose and robot’s current proprioceptive state $s_{t}$ as input, and outputs actual motor commands $a_{t}$ for the robot, where $s_{t}$ is encoded with a proprioception encoder $\mathbb{E}_{\text{prop}}$. We train low-level policy and reference motion encoder using PPO [45] in an end-to-end manner. Additionally, we introduce a motion embedding for the critic ($V_{low}$ in Figure 3) to distinguish diverse reference motions. Episodes initiate with randomized starting time steps from the dataset to avert overfitting and conclude when the root pose tracking error escalates beyond an acceptable range or the episode length is beyond the maximum length.
III-C Imitation Reward for Functionality and Style
@@
Learning Skill Functionality: To mirror the functionality of the reference motion, we translate the root pose discrepancy between agent trajectories and reference motion into a reward. The functionality reward $r_{\text{func}}$ is subdivided into tracking rewards for robot root position $r_{\text{func}}^{\text{pos}}$ and orientation $r_{\text{func}}^{\text{ori}}$. Recognizing the distinct importance of vertical movement in agile tasks, the root position tracking is further split into rewards for vertical $r_{\text{func}}^{\text{pos-z}}$ and horizontal movements $r_{\text{func}}^{\text{pos-xy}}$.
@@
The specific formulation of our functionality rewards is provided as follows, which is similar to previous work[40].
$r_{\text{func}}^{\text{ori}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-10\left\lVert\hat{\mathbf{q}}_{t}^{\text{root}}-\mathbf{q}_{t}^{\text{root}}\right\rVert^{2}\right)$
$r_{\text{func}}^{\text{pos-xy}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-20\left\lVert\hat{\mathbf{x}}_{t}^{\text{root-xy}}-\mathbf{x}_{t}^{\text{root-xy}}\right\rVert^{2}\right)$
$r_{\text{func}}^{\text{pos-z}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-80\left\lVert\hat{\mathbf{x}}_{t}^{\text{root-z}}-\mathbf{x}_{t}^{\text{root-z}}\right\rVert^{2}\right)$
@@
where $\mathbf{q},\hat{\mathbf{q}}$ and $\mathbf{x},\hat{\mathbf{x}}$ denote the root orientation and position from both the robot and reference motion, respectively. Notably, in contrast to previous work [40], we allocate a greater emphasis on root height in our reward, crucial for mastering agile locomotion skills such as backflips and jumps.
@@
Learning Skill Style: Capturing the style of a reference motion, in addition to its functionality, expands the application of the locomotion skills by meeting criteria such as energy efficiency, robot safety, and facilitating human-robot interaction. Drawing inspiration from how humans learn - starting by emulating the broader style before focusing on intricate joint movements - our robot first mimics the broader locomotion style with an adversarial style reward and later refines its technique with a joint angle tracking reward.
@@
Adversarial Stylization Reward: To quickly capture the style of the locomotion skill, we train distinct discriminators $D_{i},\ i=1..N$ for all $N$ reference motions separately to distinguish robot transitions from the transition of the specific reference motion [41, 54] and use the output to provide high-level feedback to the agent. Specifically, our discriminator is trained with the following objective:
$\mathop{\mathrm{argmin}}_{D_{i}}$ $\mathop{\mathbb{E}}_{d_{i}^{\mathcal{M}}(s_{t},s_{t+1})}\left(D_{i}(s_{t},s_{t+1})-1\right)^{2}$
$+$ $\mathop{\mathbb{E}}_{d_{i}^{\pi}(s_{t},s_{t+1})}\left(D_{i}(s_{t},s_{t+1})+1\right)^{2}$
where $d_{i}^{\mathcal{M}}(s_{t},s_{t+1})$ and $d_{i}^{\pi}(s_{t},s_{t+1})$ denote the state transition pair distribution sampled from the dataset, and the state transition pair distribution generated by the policy for $i$th reference motion respectively.
For each reference motion, the likelihood from the discriminator is then converted to a reward with:
$r_{\text{style}}^{\text{adv}}(s_{t},s_{t+1})=1-\frac{1}{4}*\left(1-D(s_{t},s_{t+1})\right)^{2}$
@@
Initially, our adversarial stylization reward provides dense reward and enables the robot to learn a credible gait, but it can not provide more detailed instructions as the training proceeds, which leads to mode collapse and unstable training.
@@
Joint Angle Tracking Reward: On the contrary, joint angle tracking reward [39, 5] provides sparse but stable instruction for the robot to mimic the gait of reference motion. Similar to our root pose tracking reward, our joint angle tracking reward has the following formulation:
@@
$r_{\text{style}}^{\text{joint}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-5\sum_{j\in\mathrm{joints}}\left\lVert\hat{\mathbf{q}}_{t}^{j}-\mathbf{q}_{t}^{j}\right\rVert^{2}\right)$
$+\exp\left(-20\sum_{f\in\mathrm{feet}}\left\lVert\hat{\mathbf{e}}_{t}^{f}-\mathbf{e}_{t}^{f}\right\rVert^{2}\right)$
$+\exp\left(-20\sum_{f\in\mathrm{feet}}\left\lVert\hat{\mathbf{h}}_{t}^{f}-\mathbf{h}_{t}^{f}\right\rVert^{2}\right)$

### v3

@@ Abstract
The agility of animals, particularly in complex activities such as running, turning, jumping, and backflipping, stands as an exemplar for robotic system design. Transferring this suite of behaviors to legged robotic systems introduces essential inquiries: How can a robot learn multiple locomotion behaviors simultaneously? How can the robot execute these tasks with a smooth transition? How to integrate these skills for wide-range applications? This paper introduces the Versatile Instructable Motion prior (VIM) – a Reinforcement Learning framework designed to incorporate a range of agile locomotion tasks suitable for advanced robotic applications. Our framework enables legged robots to learn diverse agile low-level skills by imitating animal motions and manually designed motions. Our Functionality reward guides the robot’s ability to adopt varied skills, and our Stylization reward ensures that robot motions align with reference motions. Our evaluations of the VIM framework span both simulation and the real world. Our framework allows a robot to concurrently learn diverse agile locomotion skills using a single learning-based controller in the real world. Videos can be found on our website: https://rchalyang.github.io/VIM/
@@
Figure 1: Real-Robot Trajectory. Our robot demonstrates diverse agile locomotion skills, including running, jumping, and back-flipping in real using a single motion prior and without fine-tuning.
00footnotetext: Equal Contributions
Keywords: Legged Robots, Imitation Learning, Agile Locomotion
1 Introduction
@@
Researchers have been studying for years equipping legged robots with agility comparable to that of natural quadrupeds. Picture a golden retriever gracefully maneuvering in a park: darting, leaping over obstacles, and pursuing a thrown ball. These tasks, effortlessly performed by many animals, remain challenging for contemporary legged robots. To accomplish such tasks, robots need not only master individual agile locomotion skills like running and jumping, but also the capacity to adaptively select and configure these skills based on sensory inputs. The inherent ability of quadrupeds to smoothly execute diverse locomotion skills across varied tasks inspires our pursuit of a control system with a general locomotion motion prior that includes these skills. We introduce a novel RL framework, Versatile Instructable Motion prior (VIM) aiming to endow legged robots with a spectrum of reusable agile locomotion skills by integrating existing agile locomotion knowledge.
@@
Agile gaits[1, 2, 3] for legged robots have been sculpted using model-based or optimization methods at the price of demanding significant engineering input and precise state estimation. Imitation-based controllers are also proposed to learn from motion sequences from animals [4] or optimization methods [5]. Recent works [6, 7, 8, 9, 10, 11, 12, 13, 14] also incoporate perception for legged robots. Despite encouraging results, most of these works focus on building a single controller from scratch, even though much of the learned locomotion skills could be shared across tasks. Recent works build reconfigurable low-level motion priors [15, 16, 17, 18, 19, 20] for downstream applications, but fail to make the best use of existing skills to learn diverse locomotion skills with high agility.
@@
Figure 2: Our system learns a single instructable motion prior, from a diverse reference motion dataset.
@@
In this work, we focus on building low-level motion prior to utilize existing locomotion skills in nature and previous optimization methods, and learn multiple highly agile locomotion skills simultaneously, as in Figure 2. We utilize motion sequences to offer a consistent representation of diverse agile locomotion skills. Our motion prior extracts and assimilates a range of locomotion skills from reference motions, effectively mirroring their dynamics. These references comprise motion capture (mocap) sequences from quadrupeds, synthetic motion sequences complementing mocap data, and optimized motion trajectories. We translate varied reference motion clips into a unified latent command space, guiding the motion prior to recreate locomotion skills based on these latent commands and the robot’s state. For legged robots, a locomotion skill is the ability to produce a specific trajectory. We classify this into two aspects: Functionality and Style. Functionality involves fundamental movement objectives, like moving forward at a set speed, while Style focuses on how a robot accomplishes a task, for example, two robots could run at the same speed, but with different gait. Teaching both aspects simultaneously is challenging [21]. We use three feedback: objective performance metrics, qualitative assessments, and detailed kinematic guidance. This structured approach helps the robot master basic functional objectives before refining its locomotion gaits.
@@
By incorporating diverse reference motions and our reward design, our VIM learns diverse agile locomotion skills and makes them available for intricate downstream tasks. We evaluate our method in the simulation and real world, as Figure 1. Our method significantly outperforms baselines in terms of final performance and sample efficiency.
2 Related Work
Table 1: Comparison of Skill Learning Framework. (More Details in Appendix D)
Function Style Agility Control Multiple Diverse Reusable No Privileged Real
Tracking Tracking Skills to Learn Skills Sources Information Deployment
[…]
VIM
@@ 2 Related Work
Blind Legged Locomotion: Classical legged locomotion controllers [24, 25, 26, 27, 1, 28] based on model-based methods [29, 30, 31, 32, 33, 34, 35] and trajectory optimization [36, 3] have shown promising results in diverse tasks with high levels of agility. Nonetheless, these methods normally come with considerable engineering work for the specific task, high computation requirements during deployment, or fragility to complex dynamics. Learning-based methods [6, 37, 13, 38, 39, 40, 41] controllers are proposed to offer robust and lightweight controllers for deployment at the cost of offline computation. Peng et al [42] developed a controller producing non-agile life-like gaits by imitating animals. Though previous works offer robust or agile locomotion controllers across complex environments, these works focus on finishing a single task at a time without reusing previous experience. Peng et al [22] leverage reference motions as prior knowledge when directly addressing specific tasks. Li et al [23] obtain an agile locomotion skill from a single partial rough demonstration including only robot root trajectory. Smith et al. [43] utilize existing locomotion skills to solve specific downstream tasks. Hoeller et al [44] built multiple individual locomotion skills and utilized them for agile navigation. Vollenweider et al. [45] utilize multiple AMP [22] to develop a controller to solve a fixed task set. In this paper, our motion prior captures diverse agile locomotion skills from reference motions including mocap trajectories and trajectories generated by trajectory optimization, and provides them for intricate future downstream tasks.
@@
Motion Priors: Due to the low sample efficiency and considerable effort required for reward engineering of RL, low-level skill pre-training has drawn growing attention. Singh et al [15] utilize a flow-based model to build an actionable motion prior with motion sequences generated by scripts. More recent works [16, 17, 18, 19, 46, 20, 47] focus on building low-level motion prior for downstream tasks but fail to include diverse highly agile locomotion skills. Luo et al [48, 49] developed unified motion prior for simulated humanoid robot. Peng et al [18] develop a simulation-based low-level motion prior entirely through unsupervised methods, yet they do not assure the acquisition of agile skills from the reference motion dataset (Additional discussion in Appendix C). In this work, we build motion prior with reference motions consisting of mocap sequences, synthesized motion sequences, and trajectories from optimization methods and learn multiple highly agile locomotion skills with a single controller. Comprehensive comparison is provided in Table 1.
3 Learn Versatile Instructable Motion Prior (VIM)
Building Versatile Instructable Motion prior (VIM), as shown in Figure 3, involves: constructing a reference motion dataset, and training the motion prior with an imitation-based reward system.
@@ 3 Learn Versatile Instructable Motion Prior (VIM)
Reference motion dataset: Our dataset includes $N$ reference motions for various locomotion skills such as cantering, turning, backflips, and jumps. Reference motions are from: (a) Mocap data [50] of quadrupeds; (b) synthesized motions from a generative model[50] to enhance diversity; (c) motions from trajectory optimization methods. While mocap and synthesized motions provide extensive data, not all are feasible for the robot. Thus, trajectory-optimized motions are included for complex moves like jumps and backflips. The detailed motion list is in the Appendix A. To address differences between animals and our robot, we retarget mocap and synthesized sequences as per Peng et al. [4]. Each trajectory is noted as $(s^{\text{ref}}_{0},\cdots,s^{\text{ref}}_{T})$, where $s^{\text{ref}}_{i}$ is the reference robot state at $i$th timestep, and $T$ is the length of the reference motion. We denote the dataset as $\mathcal{D}=\{(s^{\text{ref}}_{0},\cdots,s^{\text{ref}}_{T})_{i}\}_{i=1}^{N}$. Each frame includes the robot’s pose, velocity, foot position, foot height, and joint angle and velocity without motor commands. Privileged information like robot pose and velocity is used only in simulation, and policies do not require it in the real world.
3.1 Motion Prior Structure
@@
Figure 3: VIM and Reward: Our reference motion encoder maps reference motions into latent skill space and low-level policy output motor command. $V_{low}$ is the low-level critic for RL training. Our reward encourages the robot to track the root trajectory and the joint motion of the reference motion.
@@
Reference motion encoder: Our reference motion encoder $\mathbb{E}_{\text{ref}}(\cdot)$ maps segments of reference motion to latent commands in a latent skill space that outlines the robot’s prospective movement. These segments are expressed as $\hat{s}^{\text{ref}}_{t}=\{s^{\text{ref}}_{t+1},s^{\text{ref}}_{t+2},s^{\text{ref}}_{t+5},s^{\text{ref}}_{t+10},s^{\text{ref}}_{t+30}\}$. Specifically, we choose $s^{\text{ref}}_{t+1},s^{\text{ref}}_{t+2},s^{\text{ref}}_{t+5}$ to provide immediate desired future motion, $s^{\text{ref}}_{t+10},s^{\text{ref}}_{t+30}$ to provide desired motion over a longer time-span. We model the latent command as a Gaussian distribution $\mathcal{N}(\mathbb{E}_{\text{ref}}^{\mu}(\hat{s}^{\text{ref}}_{t}),\mathbb{E}_{\text{ref}}^{\sigma}(\hat{s}^{\text{ref}}_{t}))$ from which we draw a sample at each interval to guide the low-level policy. To maintain a temporal-consistent latent skill space, our training integrates an information bottleneck [51, 52] objective $L_{\text{AR}}$, where the prior follows an auto-regressive model [53]. Specifically, given the sampled latent command for the previous time step $z_{t-1}$, we minimize the KL divergence between the current latent Gaussian distribution and a Gaussian prior parameterized by $z_{t-1}$, $L_{\text{AR}}(\hat{s}^{\text{ref}}_{t},z_{t-1})=\beta\text{KL}\left(\mathcal{N}(\mu_{t},\sigma_{t}^{2})\parallel\mathcal{N}(\alpha z_{t-1},(1-\alpha^{2})I)\right)$ where $\alpha=0.95$ is the scalar controlling the effect of correlation, $\beta$ is the coefficient balancing regularization.
Low-level policy training: Our low-level policy $\pi_{\text{low}}$ takes latent command $z_{t}$ and robot’s proprioceptive state $s_{t}$ and outputs motor commands $a_{t}$ for the robot, where $s_{t}$ is encoded with a proprioception encoder $\mathbb{E}_{\text{prop}}$. We train low-level policy and reference motion encoder using PPO [54] in an end-to-end manner. We introduce learnable motion embeddings for the critic ($V_{low}$ in Figure 3) to distinguish reference motions. Episodes initiate with random frames of the dataset and terminate when the root pose tracking error is too large or the episode length is beyond the maximum length.
3.2 Imitation Reward for Functionality and Style
@@
Learning Skill Functionality: To mirror the functionality of the reference motion, we translate the root pose discrepancy between agent trajectories and reference motion into a reward. The functionality reward $r_{\text{func}}$ includes tracking rewards for robot root position $r_{\text{func}}^{\text{pos}}$ and orientation $r_{\text{func}}^{\text{ori}}$. Recognizing the distinct importance of vertical movement in agile tasks, the root position tracking is further split into rewards for vertical $r_{\text{func}}^{\text{pos-z}}$ and horizontal movements $r_{\text{func}}^{\text{pos-xy}}$.
@@
The formulation of our functionality rewards is provided as follows, similar to previous work[4]: $r_{\text{func}}^{\text{ori}}(s_{t},\hat{s}^{\text{ref}}_{t})=\exp\left(-10\left\lVert\hat{\mathbf{q}}_{t}^{\text{root}}-\mathbf{q}_{t}^{\text{root}}\right\rVert^{2}\right)$, $r_{\text{func}}^{\text{pos-xy}}(s_{t},\hat{s}^{\text{ref}}_{t})=\exp\left(-20\left\lVert\hat{\mathbf{x}}_{t}^{\text{root-xy}}-\mathbf{x}_{t}^{\text{root-xy}}\right\rVert^{2}\right)$, $r_{\text{func}}^{\text{pos-z}}(s_{t},\hat{s}^{\text{ref}}_{t})=\exp\left(-80\left\lVert\hat{\mathbf{x}}_{t}^{\text{root-z}}-\mathbf{x}_{t}^{\text{root-z}}\right\rVert^{2}\right)$ where $\mathbf{q},\hat{\mathbf{q}}$ and $\mathbf{x},\hat{\mathbf{x}}$ are the root orientation and position from the robot and reference motion respectively. Unlike previous work [4], we emphasize root height in our reward, crucial for mastering agile locomotion skills such as backflips and jumps.
@@
Learning Skill Style: Capturing the style of a reference motion, in addition to its functionality, expands the application of the locomotion skills by meeting criteria such as energy efficiency, and robot safety. Drawing inspiration from how humans learn [55, 56] - starting by emulating the broader style before focusing on intricate joint movements - our robot first mimics the broader locomotion style with an adversarial style reward and later refines its technique with a joint angle tracking reward.
@@
Adversarial Stylization Reward: We train discriminators $D_{i},\ i=1..N$ for all $N$ reference motions separately to distinguish robot transitions from the transition of reference motion [22, 45] and use the output to provide high-level feedback to the agent. Our discriminator is trained with:
$\mathop{\mathrm{argmin}}_{D_{i}}$ $\mathop{\mathbb{E}}_{d_{i}^{\mathcal{M}}(s_{t},s_{t+1})}\left(D_{i}(s_{t},s_{t+1})-1\right)^{2}+\mathop{\mathbb{E}}_{d_{i}^{\pi}(s_{t},s_{t+1})}\left(D_{i}(s_{t},s_{t+1})+1\right)^{2}$
@@
where $d_{i}^{\mathcal{M}}(s_{t},s_{t+1})$ and $d_{i}^{\pi}(s_{t},s_{t+1})$ denote the state transition pair distribution of the dataset, and the state transition pair distribution generated by the policy for $i$th reference motion respectively. For each reference motion, the likelihood from the discriminator is then converted to a reward with: $r_{\text{style}}^{\text{adv}}(s_{t},s_{t+1})=1-\frac{1}{4}*\left(1-D(s_{t},s_{t+1})\right)^{2}$ Initially, our adversarial stylization reward provides dense reward and enables the robot to learn a credible gait, but it can not provide more detailed instructions as the training proceeds, which leads to mode collapse and unstable training.
@@
Joint Angle Tracking Reward: On the contrary, joint angle tracking reward [57, 17] provides stable instruction for the robot to mimic the gait of reference motion, while the reward signal is small in scale during the initial training stage when the joint angle is away from joint target. Similar to our root pose tracking reward, our joint angle tracking reward has the following formulation:
@@
$r_{\text{style}}^{\text{joint}}(s_{t},\hat{s}^{\text{ref}}_{t})$ $=\exp\left(-5\sum_{j\in\mathrm{joints}}\left\lVert\hat{\mathbf{q}}_{t}^{j}-\mathbf{q}_{t}^{j}\right\rVert^{2}\right)+\exp\left(-20\sum_{f\in\mathrm{feet}}\left\lVert\hat{\mathbf{e}}_{t}^{f}-\mathbf{e}_{t}^{f}\right\rVert^{2}\right)+\exp\left(-20\sum_{f\in\mathrm{feet}}\left\lVert\hat{\mathbf{h}}_{t}^{f}-\mathbf{h}_{t}^{f}\right\rVert^{2}\right)$